Two wires both carry the same current I. The moving charged particles in both wires are electrons, but the con?

Two wires both carry the same current I. The moving charged particles in both wires are electrons, but the concentration of moving electrons in wire #1 is double the concentration of moving electrons in wire #2.

nder what circumstances will the drift speed of electrons be the same in both wires?

if the diameter of wire #2 is 4 times the diameter of wire #1
if the diameter of wire #2 is \sqrt 2 times the diameter of wire #1
if the diameter of wire #2 is 1/\sqrt 2 the diameter of wire #1
if the diameter of wire #2 is 2 times the diameter of wire #1
if the diameter of wire #2 is 1/2 the diameter of wire #1
if the diameter of wire #2 is 1/4 the diameter of wire #1
Answer #1

for wire1, inital DRIFT velocity is Vd = i/2neA

for wire 2, Vd = i/neA

here we must get Vd to be same for certain conditions

i,e i1/(2neA) = i2/(neA)

here i1=,i2

so 1/2A1 = 1/A2

A2 = 2A1

but A = pid^2/4

pid1^2/4 = 2* Pi d2^2/4

d2^2 = 2* d1^2

so d2 = sqrt2 times of d1

so
if the diameter of wire #2 is \sqrt 2 times the diameter of wire #1 is the answert
Answer #2

initial drift velocity Vd =i/2neA

so to get this we hav eto see the folloeing codtion to be fullfilled
if the diameter of wire #2 is \sqrt 2 times the diameter of wire #1